1)   With regard to achievement of stability of equilibrium in an isolated market assume that the market for string beans is found to have a lagged supply response such that the demand and supply function may be written as



Dt = aPt + b


St = AP t-1 + B



Required: Derive the conditions for dynamic stability of equilibrium



1
Expert's answer
2021-10-05T13:52:24-0400

Given:

"Dt = aPt + b\\ St = APt-1 + BDt=aPt+b"

At the equilibrium level,

"Dt = St\\ aPt + b = APt-1 + B .....(1)Dt=St"

If the equilibrium is stable it implies that the price level in all the periods is equal to the stable prices, i.e P*.

"=>aP \n\u2217\n +b=AP \n\u2217\n +B.....(2)"

Subtracting 1 from 2 we get,

"a (P^* - Pt) = A(P^* - Pt-1)\\ aP1 = AP2a(P 
∗
 −Pt)=A(P 
∗
 −Pt−1)"

Where P1 = Deviation of Pt from P*

P2 = Deviation of Pt-1 from P*

"\frac{P_1}{P_2} = \frac{A}{a}" , where "\frac{A}{a}" is a constant denoted by C

"\\frac{P_1}{P_2} =C \nP \n2\n\u200b\n \nP \n1\n\u200b\n \n\u200b\n =C"